OpenStudy (anonymous):

A bus travels 440 miles in 8 hours. Which unit rate describes the speed of the bus?

OpenStudy (jeffyblood):

Are there any choices?

OpenStudy (freckles):

speed =distance/time

OpenStudy (freckles):

so for example a bus travels a miles in t hours the speed of the bus is given by: \[\frac{a \text{ miles }}{ t \text{ hours } } \\ \text{ to write as a unit rate just divide }a \text{ by }t \\ \frac{\frac{a}{t} \text{ miles }}{1 hours}\]